OpenStudy (anonymous):

The doubling period of a baterial population is 20 minutes. At time t = 80 minutes, the baterial population was 80000. With t representing minutes, the formula for the population is p(t)=A e^{kt}. 1. k = . 2. The initial population at time t = 0 is . 3. The size of the baterial population after 5 hours is .

OpenStudy (anonymous):

why use this when you know the doubling time?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

its the last problem i dont understand thank you

OpenStudy (anonymous):

at \(t=80\) the population was \(8000\) the doubling time is \(20\) minutes, so at \(t=60\) it was \(4000\) and at \(t=40\) it was \(2000\) at \(t=20\) it was \(1000\) and at \(t=0\) it was \(500\)

OpenStudy (anonymous):

that answers problem #2

OpenStudy (anonymous):

i got that already and its 5000 not 500

OpenStudy (anonymous):

i just dont know 3 and 1

OpenStudy (anonymous):

3. The size of the baterial population after 5 hours is . you start with \(5000\) and 5 hours has \(5\times 3=15\) 20 minute periods (15 doublings ) so the population will be an astronomical\[5000\times 2^{15}\]

OpenStudy (anonymous):

and how do you find K?

OpenStudy (anonymous):

anyone?

OpenStudy (wolf1728):

zencvbnmjkl - what you actually need is the FORMULA right?

OpenStudy (anonymous):

yea i just need to find K

OpenStudy (wolf1728):

What is 'A' supposed to be?

OpenStudy (anonymous):

its 2=e^20k

OpenStudy (wolf1728):

at 80 minutes population = 80,000 p(t)=A e^(kt) population = A*2.7181828^(k*t) 80,000 = A * 2.7181828^(k*80) Seems we need to know 'A' you said it is its 2=e^20k but that's not a value - that's an equation What is te value of 'A'?

OpenStudy (wolf1728):

I'm guessing that 'A' is the beginning amount of bacteria right?

OpenStudy (anonymous):

i got the answer k=.03...

OpenStudy (anonymous):

i forgot the rest but thank you very much for your time

OpenStudy (anonymous):

someone gave me the answer in my class

OpenStudy (wolf1728):

After 80 minutes, population = 80,000 It doubles every 20 minutes so we have 4 doublings occurring. 2^4 = 16 So, the starting population was 80,000 / 16 = 5,000

OpenStudy (anonymous):

yea i got that but thanks for your help ma dude

OpenStudy (wolf1728):

I can show you how to solve for 'k' if you want

OpenStudy (anonymous):

do you know how?

OpenStudy (wolf1728):

Sure do

OpenStudy (anonymous):

i know it involves ln

OpenStudy (anonymous):

its ln2/20=.0346

OpenStudy (wolf1728):

yes ln means natural log - you are dealing with natural logs once 'e' is brought into an equation

OpenStudy (wolf1728):

80,000 = A * 2.7181828^(k*80) We know A = 5,000 80,000 = 5,000 * 2.7181828^(k*80) 80,000 / 5,000 = 2.7181828^(k*80) 16,000 = 2.718281828^(k*80) taking natural logs of both sides ln (16,000) = k*80 9.6803440012 / 80 = k k = 0.1210043

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

okay that's about 4 times bigger than the answer you gave me I can check the answer if you want

OpenStudy (wolf1728):

80,000 = 5,000 * 2.7181828^(k*80) We'll assume k = 0.1210043 80,000 = 5,000 * 2.7181828^(0.1210043*80) 80,000 = 5,000 * 16,000 80,000 = 80,000 Successful !!!